Find the length of the following segment created by these pairs of endpoints: (1, 1), (-3, 6)
Answer by jim_thompson5910(28536)

Note: the length of the segment is the same as the distance between the two points.

d = sqrt((x2-x1)^2+(y2-y1)^2)


d = sqrt((-3-1)^2+( 6- 1)^2)


d = sqrt((-4)^2+(5)^2)


d = sqrt(16+25)


d = sqrt(41)


So exact distance between the two points is sqrt(41) units.

Therefore, the approximate length of the segment is about 6.403 units (use a calculator to evaluate the square root of 41)